Sherlock Hemlock thought of two numbers. Their sum was 20 and the sum of their squares was 300. What was their product?
Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12844)

Their sum was 20 and the sum of their squares was 300.
What was their product?
x^2+y^2=300
x+y=20 square both sides of this equation.
x^2+2xy^y^2=400 now subtract the first equation
-(x^2+y^2=300)
--------------------
2xy=100
xy=100/2
xy=50 ans.

Answer by solver91311(24713)







Substitute:













Verification of the last two steps left as an exercise for the student.

Product is then the difference of 2 squares:




John